### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) is proposed as a cost-effective way to transfer pre-trained models to downstream tasks, avoiding the high cost of updating entire large-scale pre-trained models (LPMs). In this work, we present Fine-grained Prompt Tuning (FPT), a novel PEFT method for medical image classification. FPT significantly reduces memory consumption compared to other PEFT methods, especially in high-resolution contexts. To achieve this, we first freeze the weights of the LPM and construct a learnable lightweight side network. The frozen LPM takes high-resolution images as input to extract fine-grained features, while the side network is fed low-resolution images to reduce memory usage. To allow the side network to access pre-trained knowledge, we introduce fine-grained prompts that summarize information from the LPM through a fusion module. Important tokens selection and preloading techniques are employed to further reduce training cost and memory requirements. We evaluate FPT on four medical datasets with varying sizes, modalities, and complexities. Experimental results demonstrate that FPT achieves comparable performance to fine-tuning the entire LPM while using only 1.8% of the learnable parameters and 13% of the memory costs of an encoder ViT-B model with a 512 $\times$ 512 input resolution.
@@
Parameter-efficient fine-tuning Memory-efficient fine-tuning High-resolution Medical image classification.
@@ 1 Introduction
By utilizing the technique of fine-tuning [26], pre-trained models can be effectively adapted to specific downstream tasks by initializing the task-specific model with the weights from the pre-trained model. Recently, the remarkable achievements of large-scale pre-trained models (LPMs) [5, 14, 19, 20] further underscore the importance of this technique. However, as the model size grows rapidly, fine-tuning the parameters of an entire LPM has become very costly. To address this challenge, the concept of parameter-efficient fine-tuning (PEFT) [9, 10, 13, 23, 27, 29] has emerged, offering a strategic approach for transferring pre-trained models. PEFT involves selectively updating a small subset of pre-trained parameters or introducing a modest number of additional parameters specific to new tasks, while keeping the majority of pre-trained parameters frozen.
PEFT has successfully established its effectiveness in computer vision, but the field of medical image has not fully benefited from such advances yet [7], mainly because of the domain gap between them. In natural images, the objects of interest typically occupy a large portion of the image and exhibit distinct characteristics. In contrast, diagnostic cues in medical images often occupy a small portion and are distributed throughout the entire image [12]. Providing such fine-grained information often requires the use of high-resolution input images [2, 11]. However, as shown in Fig. 2, this preference for high-resolution images comes at the cost of increased GPU memory consumption and training expenses.
In this work, we propose a novel parameter and memory efficient fine-tuning method, namely Fine-grained Prompt Tuning (FPT). FPT aims to enhance the effectiveness of PEFT specifically for medical images in high-resolution contexts by addressing two main challenges:
1) How to efficiently extract fine-grained information from high-resolution images? Existing PEFT methods involve training a subset of parameters within the large-scale pre-trained model (LPM) [9, 10]. In contrast, FPT utilizes a lightweight additive network inspired by the concept of a side network [23, 29]. This learnable side network is introduced outside the LPM, eliminating the need for back-propagation through the LPM. However, training the side network can still be computationally expensive with high-resolution input images due to the long input sequence. FPT addresses this concern by strategically introducing asymmetric input resolution and employing important token selection to significantly reduce the length of the input sequence for the learnable network.
@@
    We present a novel PEFT method, namely Fine-grained Prompt Tuning (FPT), for medical image classification in high-resolution contexts. Asymmetric input and important token selection are proposed to improve memory efficiency. Fine-grained prompts and fine-grained fusion module are introduced to adapt pre-trained knowledge effectively and efficiently. Our code is available online.
@@
    To the best of our knowledge, this is the first work to enhance the efficiency of fine-tuning in high-resolution settings, which is particularly significant in the field of medical image analysis.
@@
    We introduce a new metric to evaluate the trade-off between performance and memory efficiency for fine-tuning methods.
@@ 2.1 Side Tuning
As illustrated in part (a) of Fig. 3, the FPT framework consists of two networks: a frozen LPM $M$ and a learnable side network $S$. Unlike other PEFT methods that introduce additional learnable parameters within the LPM, in our approach, the entire LPM remains frozen while the side network is kept learnable and separate. We adopt a lightweight architecture design for the side network, which is a scaled-down variant of the LPM. The hidden dimensions of the side network are $1/k$ times that of the LPM, where $k$ represents a reduction factor. To leverage the pre-trained knowledge from the LPM, the side network reuses the intermediate features at each layer of the LPM. Specifically, given two models $M$ and $S$ with $L$ layers, parameterized by $\theta_{M}$ and $\theta_{S}$ respectively, the intermediate activation $z_{S,\mathrm{I}}^{l}$ of layer $l$ is obtained as follows:
@@ 2.3 Fine-grained Prompts and Fusion Module
In this section, we introduce our proposed fusion module $\mathcal{F}$. As shown in the part (b) of Fig. 3, we utilize the cross-attention mechanism [1] inside the FFM to fuse features from the LPM to the side network. In the context of cross-attention, we reuse the key and value from the self-attention layer of the pre-trained model $M$. Regarding the query, one approach is to directly reuse the query from the side network $S$. However, the cross-attention map can be large if the input sequence is long, leading to increased memory consumption. Therefore, we introduce a small set of learnable embeddings $z_{p}$, namely fine-grained prompts, into each layer of the fusion modules as the query. Unlike prompt tuning [13, 16], which directly uses prompts as part of the input sequence, fine-grained prompts serve as a bridge linking the frozen LPM and the side network. These prompts are concatenated with the intermediate sequence of the side network to join the forward propagation after fusing pre-trained features from the LPM. Specifically, the fusion module is processed as follows:
$\mathcal{F}(z_{S},z_{M})=f_{out}([z_{S},\text{CA}(z_{p},z_{M})]),$ (3)
@@ 3.2.1 Experiment setup
In this study, we utilize a popular variant of Vision Transformers (ViT) [6], specifically ViT-B(ase), which is pre-trained on ImageNet-21K [21]. All methods are fine-tuned for 20 epochs with a mini-batch size of 16. We use the AdamW optimizer [18] with cross-entropy as the loss function for all datasets. To ensure fair comparisons, we conduct a grid search of hyper-parameters for all methods. All methods are run at a resolution of $512\times 512$. In the context of FPT, the high input resolution for the LPM remains $512\times 512$, while the low input resolution for the side network is set to $224\times 224$. We set the reduction factor $k$ to 8 and use 16 fine-grained prompts with the same hidden dimension as $S$. For important token selection, we retain the top 20% of important tokens.
@@ 3.2.2 Evaluation metric
Table 1: Comparison results with state-of-the-art PEFT methods across four evaluation datasets. ‘Params.’ denotes the ratio of learnable parameters to the total number of parameters. ‘Mem.’ denotes the memory usage (MB). The best and second-best results are bold and underlined, respectively.
@@ 3.3 Comparisons with State-of-the-art
We compare FPT against full fine-tuning, linear probing, and state-of-the-art (SOTA) PEFT approaches. In full fine-tuning, all parameters of the LPM are made learnable during training on the downstream tasks. Linear probing involves solely training the new task-specific head on top of the LPM. Generally, full fine-tuning often represents the upper performance bound for fine-tuning, while linear probing represents the lower bound. We also compare FPT against five other popular SOTA PEFT methods.
The performance and efficiency of PEFT methods are tabulated in Table 1. It can be observed that fine-tuning the entire LPM (full fine-tuning), as the upper bound, achieves the best average AUC of 93.96%, requiring 24,116MB of training GPU memory with a batch size of 16. Although other PEFT methods improve fine-tuning efficiency by reducing the number of learnable parameters, they are unable to reduce the overhead brought by the high input resolution; all compared PEFT methods require at least 20, [^0] MB of memory, and even linear probing requires 4,364MB of memory. In contrast, FPT achieves the second-best AUC of 92.26%, while significantly reducing the memory requirement to only 3,182MB (13% of full fine-tuning). Moreover, with a separate network, FPT utilizes only 1.81% learnable parameters compared to full fine-tuning. Therefore, in terms of efficiency, FPT achieves the best PPE and PME, presenting the best trade-off between performance and efficiency. This parameter and memory efficiency positions FPT as a practical and feasible choice for leveraging LPMs in high-resolution contexts.
@@ 4 Conclusion
In this paper, we introduce a novel PEFT method, namely Fine-grained Prompt Tuning (FPT), for medical image classification. FPT significantly reduces memory requirements, particularly in the high-resolution context that commonly used in medical image analysis. To address the challenge of high memory requirement, we first adopt the design of side tuning and enhance it with an asymmetric input strategy. We then introduce fine-grained prompts and the fine-grained fusion module to allow effective adaptation of pre-trained knowledge from an out-of-domain LPM that takes images of a different scale as input. To further reduce memory requirement, important token selection and the preloading of pre-trained features are applied. By integrating these components, our PEFT method achieves superior performance across four medical datasets while maintaining the best parameter and memory efficiency.

### v3

@@ Abstract
Parameter-efficient transfer learning (PETL) is proposed as a cost-effective way to transfer pre-trained models to downstream tasks, avoiding the high cost of updating entire large-scale pre-trained models (LPMs). In this work, we present Fine-grained Prompt Tuning (FPT), a novel PETL method for medical image classification. FPT significantly reduces memory consumption compared to other PETL methods, especially in high-resolution input contexts. To achieve this, we first freeze the weights of the LPM and construct a learnable lightweight side network. The frozen LPM takes high-resolution images as input to extract fine-grained features, while the side network is fed low-resolution images to reduce memory usage. To allow the side network to access pre-trained knowledge, we introduce fine-grained prompts that summarize information from the LPM through a fusion module. Important tokens selection and preloading techniques are employed to further reduce training cost and memory requirements. We evaluate FPT on four medical datasets with varying sizes, modalities, and complexities. Experimental results demonstrate that FPT achieves comparable performance to fine-tuning the entire LPM while using only 1.8% of the learnable parameters and 13% of the memory costs of an encoder ViT-B model with a 512 $\times$ 512 input resolution.
@@
Parameter-efficient transfer learning Memory-efficient transfer learning High-resolution medical image classification.
@@ 1 Introduction
By utilizing the technique of fine-tuning [26], pre-trained models can be effectively adapted to specific downstream tasks by initializing the task-specific model with the weights from the pre-trained model. Recently, the remarkable achievements of large-scale pre-trained models (LPMs) [5, 14, 19, 20] further underscore the importance of this technique. However, as the model size grows rapidly, fine-tuning the parameters of an entire LPM has become very costly. To address this challenge, the concept of parameter-efficient transfer learning (PETL) [9, 10, 13, 23, 27, 29] has emerged, offering a strategic approach for transferring pre-trained models. PETL involves selectively updating a small subset of pre-trained parameters or introducing a modest number of additional parameters specific to new tasks, while keeping the majority of pre-trained parameters frozen.
PETL has successfully established its effectiveness in computer vision, but the field of medical image has not fully benefited from such advances yet [7], mainly because of the domain gap between them. In natural images, the objects of interest typically occupy a large portion of the image and exhibit distinct characteristics. In contrast, diagnostic cues in medical images often occupy a small portion and are distributed throughout the entire image [12]. Providing such fine-grained information often requires the use of high-resolution input images [2, 11, 30]. However, as shown in Fig. 2, this preference for high-resolution images comes at the cost of increased GPU memory consumption and training expenses.
In this work, we propose a novel parameter and memory efficient transfer learning method, namely Fine-grained Prompt Tuning (FPT). FPT aims to enhance the effectiveness of PETL specifically for medical images in high-resolution input contexts by addressing two main challenges:
1) How to efficiently extract fine-grained information from high-resolution images? Existing PETL methods involve training a subset of parameters within the large-scale pre-trained model (LPM) [9, 10]. In contrast, FPT utilizes a lightweight additive network inspired by the concept of a side network [23, 29]. This learnable side network is introduced outside the LPM, eliminating the need for back-propagation through the LPM. However, training the side network can still be computationally expensive with high-resolution input images due to the long input sequence. FPT addresses this concern by strategically introducing asymmetric input resolution and employing important token selection to significantly reduce the length of the input sequence for the learnable network.
@@
    We present a novel PETL method, namely Fine-grained Prompt Tuning (FPT), for medical image classification in high-resolution contexts. Asymmetric input and important token selection are proposed to improve memory efficiency. Fine-grained prompts and fine-grained fusion module are introduced to adapt pre-trained knowledge effectively and efficiently. Our code is available online.
@@
    To the best of our knowledge, this is the first work to enhance the efficiency of PETL in high-resolution input settings, which is particularly significant in the field of medical image analysis.
@@
    We introduce a new metric to evaluate the trade-off between performance and memory efficiency for transfer learning methods.
@@ 2.1 Side Tuning
As illustrated in part (a) of Fig. 3, the FPT framework consists of two networks: a frozen LPM $M$ and a learnable side network $S$. Unlike other PETL methods that introduce additional learnable parameters within the LPM, in our approach, the entire LPM remains frozen while the side network is kept learnable and separate. We adopt a lightweight architecture design for the side network, which is a scaled-down variant of the LPM. The hidden dimensions of the side network are $1/k$ times that of the LPM, where $k$ represents a reduction factor. To leverage the pre-trained knowledge from the LPM, the side network reuses the intermediate features at each layer of the LPM. Specifically, given two models $M$ and $S$ with $L$ layers, parameterized by $\theta_{M}$ and $\theta_{S}$ respectively, the intermediate activation $z_{S,\mathrm{I}}^{l}$ of layer $l$ is obtained as follows:
@@ 2.3 Fine-grained Prompts and Fusion Module
In this section, we introduce our proposed fusion module $\mathcal{F}$. As shown in the part (b) of Fig. 3, we utilize the cross-attention mechanism [1] inside the FFM to fuse features from the LPM to the side network. In the context of cross-attention, we reuse the key and value from the self-attention layer of the pre-trained model $M$. Regarding the query, one approach is to directly reuse the query from the side network $S$. However, the cross-attention map can be large if the input sequence is long, leading to increased memory consumption. Therefore, we introduce a small set of learnable embeddings $z_{p}$, namely fine-grained prompts, into each layer of the fusion modules as the query. Unlike prompt tuning [13, 16], which directly uses prompts as part of the input sequence, fine-grained prompts serve as a bridge linking the frozen LPM and the side network. These prompts are concatenated with the intermediate sequence of the side network to join the forward propagation after fusing pre-trained features from the LPM. They are then removed after the layer’s forward processing. Specifically, the fusion module is processed as follows:
$\mathcal{F}(z_{S},z_{M})=[z_{S},f_{out}(\text{CA}(z_{p},z_{M}))],$ (3)
@@ 3.2.1 Experiment setup
In this study, we utilize a popular variant of Vision Transformers (ViT) [6], specifically ViT-B(ase), which is pre-trained on ImageNet-21K [21]. All methods are fine-tuned for 20 epochs with a mini-batch size of 16. We use the AdamW optimizer [18] with cross-entropy as the loss function for all datasets. To ensure fair comparisons, we conduct a grid search of hyper-parameters for all methods. All methods are run at a resolution of $512\times 512$. In the context of FPT, the high input resolution for the LPM remains $512\times 512$, while the low input resolution for the side network is set to $224\times 224$. All methods were trained employing the same data augmentations as those for low-resolution inputs in FPT. We set the reduction factor $k$ to 8 and use 16 fine-grained prompts with the same hidden dimension as $S$. For important token selection, we retain the top 20% of important tokens.
@@ 3.2.2 Evaluation metric
Table 1: Comparison results with state-of-the-art PETL methods across four evaluation datasets. ‘Params.’ denotes the ratio of learnable parameters to the total number of parameters. ‘Mem.’ denotes the memory usage (MB). The best and second-best results are bold and underlined, respectively.
@@ 3.3 Comparisons with State-of-the-art
We compare FPT against full fine-tuning, linear probing, and state-of-the-art (SOTA) PETL approaches. In full fine-tuning, all parameters of the LPM are made learnable during training on the downstream tasks. Linear probing involves solely training the new task-specific head on top of the LPM. Generally, full fine-tuning often represents the upper performance bound for transfer learning, while linear probing represents the lower bound. We also compare FPT against five other popular SOTA PETL methods.
The performance and efficiency of PETL methods are tabulated in Table 1. It can be observed that fine-tuning the entire LPM (full fine-tuning), as the upper bound, achieves the best average AUC of 93.96%, requiring 24,116MB of training GPU memory with a batch size of 16. Although other PETL methods improve transfer learning efficiency by reducing the number of learnable parameters, they are unable to reduce the overhead brought by the high input resolution; all compared PETL methods require at least 20, [^0] MB of memory, and even linear probing requires 4,364MB of memory. In contrast, FPT achieves the second-best AUC of 92.26%, while significantly reducing the memory requirement to only 3,182MB (13% of full fine-tuning). Moreover, with a separate network, FPT utilizes only 1.81% learnable parameters compared to full fine-tuning. Therefore, in terms of efficiency, FPT achieves the best PPE and PME, presenting the best trade-off between performance and efficiency. This parameter and memory efficiency positions FPT as a practical and feasible choice for leveraging LPMs in high-resolution contexts.
@@ 4 Conclusion
In this paper, we introduce a novel PETL method, namely Fine-grained Prompt Tuning (FPT), for medical image classification. FPT significantly reduces memory requirements, particularly in the high-resolution context that commonly used in medical image analysis. To address the challenge of high memory requirement, we first adopt the design of side tuning and enhance it with an asymmetric input strategy. We then introduce fine-grained prompts and the fine-grained fusion module to allow effective adaptation of pre-trained knowledge from an out-of-domain LPM that takes images of a different scale as input. To further reduce memory requirement, important token selection and the preloading of pre-trained features are applied. By integrating these components, our PETL method achieves superior performance across four medical datasets while maintaining the best parameter and memory efficiency.